OpenStudy (anonymous):

Pola wants 3 oz of a 20% solution. she has only 5% and 30% available. how many oz of each should she mix to get what she wants?

OpenStudy (anonymous):

x+y =3 oz .05*x + .3*y = .2*3

OpenStudy (anonymous):

could you please solve the equations..i dont know how to....

OpenStudy (anonymous):

and why did u multiply .2 by 3

OpenStudy (anonymous):

what class are you taking?

OpenStudy (cwrw238):

Algebraic multiplied 0.2 by 3 because thats the amount of substance in 3 oz of solution (0.2 = 20%)

OpenStudy (anonymous):

math 130- intermediate algebra

OpenStudy (anonymous):

should be no problem for you to solve. very simple substitution.

OpenStudy (anonymous):

ok. thnks

OpenStudy (anonymous):

when pppl give yu best responses yur smart score goes up huh?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

i multiplied the top equation by -.3 so that i could cancel out the y value and after i solved i got x= -.21

OpenStudy (anonymous):

solve for either x or y, your choice: x+y =3 sub.s into: .05*x + .3*y = .2*3